OpenStudy (anonymous):

4x+5y=20 How do I find the y AND x intercepts of this question?

OpenStudy (anonymous):

Incomplete

OpenStudy (anonymous):

To find the y-intercept, set x=0. To find the x-intercept, set y=0.

OpenStudy (anonymous):

I understand that part, but after that I get confused...

OpenStudy (anonymous):

y-intercept, plug in 0 for x in the equation. 4*0+5y=20 x-intercept, plug in 0 for y in the equation. 4x+5*0=20 Find x for y-intercept and y for x-intercept.

OpenStudy (anonymous):

so for the y intercept I would subtract the 4y from the 20 as well as the 4x?

OpenStudy (anonymous):

4x I mean

OpenStudy (anonymous):

I will show solution for y-intercept 4*0+5y=20 5y=20 y=4 So the y-intercept is (0,4) Try the same thing for x-intercept.

OpenStudy (anonymous):

I see, thank you!

OpenStudy (anonymous):

Welcome ^^

OpenStudy (anonymous):

so if Im doing this right, the x intercept is 5?

OpenStudy (anonymous):

(5,0) Yep.